OpenStudy (anonymous):

When the quadratic equation y = (x - 4)(2x + 5) is written in standard form, what is the value of the coefficient "b"?

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

FOIL it out...

OpenStudy (anonymous):

the b coefficient is obtained by adding O + I O - outside terms I - inside terms

OpenStudy (anonymous):

y = (x - 4)(2x + 5) y = 2x² - 3x - 20 In Standard form : y = ax² + bx + c Therefore y = 2x² - 3x - 20 Answer : b = -3